Show your Solutions

3. A circuit contains four 15-ohm resistors connected in series across a 120 volt source.

   a) What is the effective resistance of the circuit?

 

 

 

   b) What current will flow in the circuit?

 

 

 

   c) What will be the voltage across each resistor


1
Expert's answer
2021-04-23T11:42:56-0400

a) "R_{ef} = 4R = 4 \times 15 = 60 \;ohm"

b) "I = \frac{V}{R_{ef}} = \frac{120}{60} = 2 \;A"

c) "V_R = IR = 2 \times 15 = 30 \;V"